OpenStudy (anonymous):

1+2+3+4+5+6+7+8+9+10=?

OpenStudy (gorv):

n(n+1)/2

OpenStudy (anonymous):

really?? -_- don't you know addition?? lol

OpenStudy (gorv):

where n is total number here n=10 10*(10+1)/2=10811/2=55

OpenStudy (anonymous):

i just want to give a medal away

OpenStudy (anonymous):

55 Use http://www.wolframalpha.com/ for math

OpenStudy (anonymous):

1+2+3 = 6 6+4 = 10 5+10 = 15 6+15 = 21 7+21 = 28 8+28 = 36 9+36 = 45 10+45 = 55 I meant to type this earlier but my computer froze, lol